OpenStudy (andriod09):

I need help with a Pythagorian Theory problem. The problem is: A^2+b^2=c^2 X^2+(x^2+7)^2=C^2 how can i find the correct answers with out an absolute number??

OpenStudy (anonymous):

what do you mean? \[x^2+(x^2+7)^2=c^2

OpenStudy (anonymous):

\[x^2+(x^2+7)^2=c^2\]

OpenStudy (anonymous):

\[x^2+x^4+14x^2+49=c^2\]

OpenStudy (anonymous):